OpenStudy (anonymous):

Solve for x & y Equation 1: ax + by = r ax + cy = s Equation 2: cx + ry = m dx - sy = n

OpenStudy (anonymous):

ax+by=r -ax-cy=-s ------------------ by-cy=r-s y(b-c)=r-s y=r-s/(b-c)

OpenStudy (anonymous):

What about X

OpenStudy (anonymous):

let plug it in ax+b(r-s)/(b-c)=r just solve that

OpenStudy (anonymous):

Better method of solving \[\left( \begin{array}{cc} a & b \\ a & c \end{array} \right)\left( \begin{array}{c} x \\ y \end{array} \right)=\left( \begin{array}{c} r \\ s \end{array} \right)\]

OpenStudy (anonymous):

c(ax + by = r ) -b(ax + cy = s) ------------------ acx + bcy = cr -abx - bcy = -bs ----------------- acx - abx = cr - bs x(ac - ab) = cr - bs x = cr - bs / ac - ab Am I Right!

OpenStudy (anonymous):

x y \[\left\{\left\{\frac{b s-c r}{a (b-c)}\right\},\left\{\frac{r-s}{b-c}\right\}\right\}\]